OpenStudy (anonymous):

simplify the complex fraction (let's see if I can translate the format so it's understandable): (1/y) -3 over (1/y^3) -9 = ??

OpenStudy (anonymous):

\[\frac{\frac{1}{y}-3}{\frac{1}{y^3}-9}\]

OpenStudy (anonymous):

like that?

OpenStudy (anonymous):

multiply top and bottom by \[y^3\] to get \[\frac{y^2-3y^3}{1-9y^3}\]

OpenStudy (anonymous):

\[(-2/y)/8y^3/9\] =9/(4Y^2)

OpenStudy (anonymous):

(1/y - 3y/y)/(y^3/y^3 - y^3/9)